OpenStudy (anonymous):

Add 3/3 + 5/6 + 4/12

OpenStudy (unklerhaukus):

fist you'll want to convert those fraction to have a common denominator what is the lowest common multiple of 3, 6 and 12?

OpenStudy (anonymous):

is it 12 ?

OpenStudy (unklerhaukus):

yes

OpenStudy (unklerhaukus):

what do you have to multiply 3 by to get 12?

OpenStudy (anonymous):

4.

OpenStudy (unklerhaukus):

ok so what does 3/3 become when you multiply the numerator and denominator by 4

OpenStudy (unklerhaukus):

3/3=(4x3)/(4x3)=

OpenStudy (anonymous):

11/6 + 4/12 = ((11 × 12) + (4 × 6)) / (6 × 12) = (132 + 24) / 72 = 156/72 = 13/6 = 2 1/6

OpenStudy (anonymous):

so thats ur answer

OpenStudy (anonymous):

if u ever need help on fractions just message me and i can help!

OpenStudy (anonymous):

anymore

OpenStudy (anonymous):

u still on???